• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name flay

    Log entry time 00:02:35 on May 8,2010

    Entry number 321614

    keyword=5/7/10 Swing Shift Summary

    16:00 Shift begins with raster tests; access to Hall (with escort) for detector work (see Halog entry).

    17:30 Request for 5uA beam for beam recovery procedure. MCC asked to give us 50uA to see if we can withstand it (we're on empty target).

    17:35 MCC test complete. Continuing beam recovery.

    18:20 Starting BPM Calibration test. Unsyncing raster.

    18:25 MCC sending in 5uA beam. spot++ looks good; going to 50uA.
    Turning Auto Gain off.

    18:55 BPM calibration runs complete. Moving back to Lead/Diamond target #3; Auto Gain turned back on.

    19:05 Continuing raster check.

    19:30 Syncing raster.

    19:40 Moving to 50uA (raster sync test).

    21:00 Beginning raster sync test.

    21:30 MCC alerted us that the Tungsten thermo-couples have been unresponsive on the strip chart since ~20:30... Going to run at ~40uA for production data later in the evening until we can solve the issue.

    22:55 Raster sync test ongoing.

    23:40 Raster sync test ongoing.